OpenStudy (anonymous):

u(y)=U[1−(y^2/b^2)] find du/dy

OpenStudy (austinl):

\(u(y)=U[1-(\dfrac{y^2}{b^2})]\) This is correct yes? And you want to take the derivative of \(U[1-(\dfrac{y^2}{b^2})]\) with respect to \(y\) correct?

OpenStudy (anonymous):

yes

OpenStudy (austinl):

Ok, let me see here.... whoa boy. I can see why you asked about this :)

OpenStudy (anonymous):

i feel like i should know this but i have been out of school for quite some time

OpenStudy (austinl):

Now I am going to try this, but I provide no guarantee as to the accuracy of my techniques. I think we need to do implicit differentiation.

OpenStudy (austinl):

\(\dfrac{du}{dy}(U[1-(\dfrac{y^2}{b^2})])\) \(=\dfrac{d}{dy}(U[1-(\dfrac{y^2}{b^2})])\)\) \(\dfrac{d}{dy}(U[1-(\dfrac{y^2}{b^2})])=\dfrac{dU(u)}{du}\dfrac{du}{dy}\) where, \(u=1-\dfrac{y^2}{b^2}\) and \(\dfrac{d}{dy}U(u)=U^{\prime}(u)\) \(=(\dfrac{d}{dy}(1-\dfrac{y^2}{b^2}))U^{\prime}(1-\dfrac{y^2}{b^2})\) Differentiate term by term and factor out any constants. \([\dfrac{d}{dy}(1)-\dfrac{\dfrac{d}{dy}(y^2)}{b^2}]U^{\prime}(1-\dfrac{y^2}{b^2})\) \(=-\dfrac{(\dfrac{d}{du}(y^2))U^{\prime}(1-\dfrac{y^2}{b^2})}{b^2}\) \(=\color{blue}{-\dfrac{(2y)U^{\prime}(1-\dfrac{y^2}{b^2})}{b^2}}\) \(\color{red}{\text{DISCLAIMER: Not entirely sure this is correct}}\)

OpenStudy (austinl):

@amistre64 Care to check my work?

OpenStudy (austinl):

@Luigi0210 Any thoughts?

OpenStudy (atlas):

Is U a function here??

OpenStudy (atlas):

@austinL

OpenStudy (anonymous):

no, for this problem it is a consant

OpenStudy (austinl):

All I know is what was given to us. And so I just put \(U^{\prime}\).

OpenStudy (atlas):

@austinL: you made an easy question look difficult :P

OpenStudy (anonymous):

that's my fault, I should have defined that better

OpenStudy (atlas):

yeah I understand that @mfred81

OpenStudy (atlas):

:)

OpenStudy (austinl):

@atlas Was I correct though? That is the important thing!

OpenStudy (atlas):

Nope I don't think so

OpenStudy (austinl):

Well gosh darn it :P

OpenStudy (atlas):

k......did you assume U as a function of u??

OpenStudy (austinl):

I tried plugging it into my calculator to slve and it just spit it back out and said nope nope nope nope. Wolfie agrees however, and also shows a different form.

OpenStudy (austinl):

solve*

OpenStudy (atlas):

u(y) = U[1-y^2/b^2] I am assuming U is a function of U and b is a constant Then du/dy = U[-2y/b^2] + [1-y^2/b^2] U' (du/dy)

OpenStudy (atlas):

du/dy(1-U' + U'(y^2/b^2)) = -2U[y/b^2]

OpenStudy (atlas):

that's the answer assuming U is a direct function of U and dU/du = U'

OpenStudy (anonymous):

thank you for your help guys! I am waaaay out of practice

OpenStudy (atlas):

so are we @mfred81